Find an equation that describes the following variation
y varies directly as z, and y = 46 when z = 460.

Answer by fractalier(6550)

Since y varies directly as z, we write that as
y = kz
Then we plug our data in to find k...
46 = k(460)
k = 46/460 = 1/10
so that
y = (1/10)z